OpenStudy (anonymous):

does anybody know the sum of 44,521,648 is ???? i'll fan and medal

OpenStudy (anonymous):

Try dividing it by 2

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

what did u get

OpenStudy (anonymous):

i got 22,543,123 or is that wrong

OpenStudy (anonymous):

ok now add 22,543,123 to 22,543,123 and tell me what you get

OpenStudy (anonymous):

22+22=44 5+5=10 4+4=8 3+3=6 1+1=2 2+2=4 3+3=6 the answer is 44,108,622,46

OpenStudy (anonymous):

22260824 that is what u get when u divide. Now, you add the numbers and you get = 44521648. B/c half of 44521648 is 22260824

OpenStudy (anonymous):

okay thanks for help

OpenStudy (anonymous):

Fan and medal or no?

OpenStudy (anonymous):

okay i'll fan and medal

OpenStudy (anonymous):

briannabear

OpenStudy (anonymous):

can you put your name in blue